OpenStudy (anonymous):

evaluate integral 0 to 1 (x)/(x^2 - 1)

OpenStudy (abb0t):

This requires u-substitution. Arre you familiar with this method?

OpenStudy (abb0t):

\[\int\limits_{0}^{1}\frac{ x }{ x^2-1 }dx\]

OpenStudy (anonymous):

no

OpenStudy (abb0t):

\(u = x^2-1\) while \(du=2x\) you don't have a 2x, so you're going ot have to divide both sides by 2. You want something that you can substitute. You have \[\int\limits_{0}^{1}\frac{ 2 }{ u }du = 2 \int\limits \frac{ 1 }{ u }du\] now clearly, you can do the integral of that, which is simply the natural log plus a constant. but you have an definite integral, so you don't need a constant. Don't for get to substitute back your original "u" which was \(u=x^2-1\) REMEMBER. some people like to change the bounds and evaluate using "u", but i find prefer to evaluate using original function, since I don't have to change anything. So byu fundamental theorem of calculus II you should know that you have \(F(b)-F(a)\) where you're bound is 0<x<1 or [0,1]

OpenStudy (anonymous):

Okay but this integral does not converge :/

OpenStudy (abb0t):

Oh, you're right. What am I doing. Lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so? should i just say that or?

OpenStudy (abb0t):

Do you know WHY it doesn't converge?first of all. :)

OpenStudy (anonymous):

ohhh lol

OpenStudy (abb0t):

As you can see, your problem is the upper limit.

OpenStudy (anonymous):

okay

OpenStudy (abb0t):

so, evaluate the integral from 0 to some variable, say "R"....and take the limit.

OpenStudy (anonymous):

umm ok

OpenStudy (abb0t):

Lol, where are you confused?

OpenStudy (anonymous):

so its convegent?

OpenStudy (anonymous):

@RadEn can u help?

OpenStudy (raden):

the integral not exist for upper interval x = 1